# Spurious Privacy Leakage in Neural Networks

### v1

@@ Abstract
Neural networks are vulnerable to privacy attacks aimed at stealing sensitive data. The risks can be amplified in a real-world scenario, particularly when models are trained on limited and biased data. In this work, we investigate the impact of spurious correlation bias on privacy vulnerability. We introduce spurious privacy leakage, a phenomenon where spurious groups are significantly more vulnerable to privacy attacks than non-spurious groups. We further show that group privacy disparity increases in tasks with simpler objectives (e.g. fewer classes) due to the persistence of spurious features. Surprisingly, we find that reducing spurious correlation using spurious robust methods does not mitigate spurious privacy leakage. This leads us to introduce a perspective on privacy disparity based on memorization, where mitigating spurious correlation does not mitigate the memorization of spurious data, and therefore, neither the privacy level. Lastly, we compare the privacy of different model architectures trained with spurious data, demonstrating that, contrary to prior works, architectural choice can affect privacy outcomes.
@@ 1 Introduction
In this work, we examine the spurious correlation bias, a statistical relationship between two variables that appears to be causal but is either caused by a third confounding variable or randomness. Such correlations are prevalent in real-world datasets, particularly in domains with limited and imbalanced group representation. To address the problem, several mitigation methods have been proposed in the literature with the objective to improve the performance of the worst-group (Sagawa et al., 2019; Nam et al., 2020; Liu et al., 2021a; Izmailov et al., 2022; Yang et al., 2023). However, their privacy implications have not been studied. On the other hand, privacy research typically focuses on unbiased datasets such as CIFAR10 or CIFAR100 (Krizhevsky, 2009; Hu et al., 2022), overlooking the privacy risks of realistic applications that make use of limited and biased real-world datasets. We address this gap by investigating the privacy of neural networks trained on real-world spurious correlated datasets using membership inference attacks (MIA), a family of privacy attacks commonly used for their simplicity and versatility (Murakonda & Shokri, 2020; Carlini et al., 2021).
@@
Contributions. We identify a phenomenon we term spurious privacy leakage, where groups with spurious correlation are significantly more vulnerable to MIA than other groups, creating a privacy disparity (Section 3.1). This phenomenon adds a fairness requirement on top to the existing privacy challenges, modeling a realistic scenario for data-sensitive applications. Prior works demonstrated how privacy auditing may naively conclude that a model satisfies the privacy requirements by evaluating on an aggregated average metric over the entire dataset (Carlini et al., 2022; Ye et al., 2022; Sablayrolles et al., 2019). We demonstrate that this also applies for realistic spurious data, where an average metric evaluation is misleading for evaluating spurious and non-spurious groups. Addressing this privacy disparity is important to bridge and advance the privacy and spurious correlation research, to characterize the risks of the model, and to improve the auditing process.
@@
We continue with further analysis on spurious privacy leakage. Prior works suggested that improving fairness can, in limited scenarios, bound privacy disparity (Kulynych et al., 2022). Given the range of methods proposed to mitigate spurious correlations (equivalent as improving fairness between different data groups) (Sagawa et al., 2019; Kirichenko et al., 2022; Liu et al., 2021a; Zhang et al., 2022), we investigate whether these mitigations can also reduce privacy leakage. To explore this, we apply robust training methods to mitigate spurious correlations and re-evaluate the privacy vulnerability. Surprisingly, we find no consistent and significant privacy improvement (Section 4.1). This observation leads us to introduce a new perspective on group privacy disparity based on memorization (Zhang et al., 2021): spurious robust training improves worst-class performance but does not reduce the memorization level of data compared to standard training (Figure 7), and thus neither can mitigate the privacy risks. Additional results demonstrate that robust training essentially learns the same features as standard training. We also assess the effectiveness of differential privacy to mitigate the spurious privacy leakage. Given a pretrained model, a stricter privacy guarantee can offer worst group privacy protection at the expense of reduced utility (Section 4.2).
@@
Lastly, while prior works mostly focus on ResNet-like architecture for privacy analysis (Carlini et al., 2022; Liu et al., 2022a), we fairly compare the performance and privacy of different model families. Our evaluation include state-of-the-art convolutional and transformer-based architectures, pretrained using both supervised and self-supervised training. Contrary to prior works showing that the best shadow architecture always matches the target one (Carlini et al., 2022; Liu et al., 2022a), our results provide counterexamples when models are trained on spurious correlated data. This highlights a key difference in privacy behavior when models are trained on biased real-world data. We release the code at https://anonymous.4open.science/r/spurious-mia-4FB5.
@@ 2 Background & Related work
where the dataset is divided into $\bm{g}$ groups. The new dataset is $\mathcal{D}=\{(\bm{x}_{i},\bm{y}_{i},\bm{g}_{i})\}$ where $\bm{g}\in{\mathcal{G}}$ is a discrete-valued label (e.g. all the combinations of birds and backgrounds or geographical area information (Koh et al., 2021)). DRO is considered an oracle method due to its explicit use of the group information for the training (Liu et al., 2021a). Additional methods in the literature suppress the spurious features by learning and assigning a different weight per sample (Liu et al., 2021a; Nam et al., 2020), by retraining the classifier head at the end of the training (Kirichenko et al., 2022; Izmailov et al., 2022; Kang et al., 2020), by group sampling (Yang et al., 2024; Idrissi et al., 2022), or using contrastive methods (Zhang et al., 2022).
@@
Privacy and safety. The intersection of privacy and ML safety topics has been extensively studied. Wang et al. (2021) focused on how pruning can mitigate privacy attacks, Shokri et al. (2021) explored the connection between privacy and explainability, Song et al. (2019) found that adversarial training can increase privacy leakage, but however, Li et al. (2024) reported contradictory findings when using a better privacy evaluation (Carlini et al., 2022). In our work, we investigate the privacy risk of real-world spurious correlated datasets, which is related to fairness where the goal is to ensure equal performance across groups. Prior works on privacy-fairness reported that subpopulations can exhibit varying levels of privacy risk (Tian et al., 2024; Zhong et al., 2022; Kulynych et al., 2022). In particular, Tian et al. (2024) showed that fairness methods can mildly mitigate MIA risks using the average metrics. Instead, we use a per-group analysis, revealing the spurious privacy leakage phenomenon with group privacy disparity between spurious and non-spurious data. Kulynych et al. (2022) and Zhong et al. (2022) also investigated privacy disparity on synthetic and tabular datasets. They hypothesize that, under certain conditions, group fairness improvements or differential privacy can be effective as privacy mitigation. In Section 4.1, we show that neither approaches can effectively address privacy disparities in real-world datasets. Similar to our setting, Yang et al. (2022) studied the privacy risks associated with a toy dataset (MNIST) using a suboptimal evaluation. In contrast, we use state-of-the-art methods to evaluate real-world datasets. Lastly, while out-of-distribution data are known to be more vulnerable to privacy attacks than in-distribution data (Carlini et al., 2022), spurious correlated data differ from out-of-distribution data, as they are by definition in-distribution. Our contributions present a set of results to resolve the conflicts in the literature while also offering new directions for future works.
@@ 3.1 Spurious privacy leakage
Experiment setup. We select the datasets that are used by the spurious correlation community (Yang et al., 2023): Waterbirds (Sagawa et al., 2019), CelebA (Liu et al., 2015), FMoW (Koh et al., 2021), MultiNLI (Williams et al., 2017), and CivilComments (Koh et al., 2021). These datasets contain real-world spurious correlations, diverse modalities, and different target complexity (see Appendix A for details). Moreover, to the best of our knowledge, we are the first to study MIA attacks on subgroups of these realistic datasets. We use the pretrained ResNet50 (He et al., 2016) on ImageNet1k and finetune using random crop and horizontal flip. For text datasets, BERT’s bert-base-uncased model (Devlin et al., 2019) is used. We perform hyperparameter optimization for each dataset using a grid search over learning rate (lr), weight decay (wd), and epochs. The grid search and its best hyperparameters are in Appendix B. We report the training and test accuracy to evaluate the performance and their difference to quantify the overfitting level. We do the same for the worst-group accuracy (WGA), which is a commonly used proxy metric to measure the mitigation success of spurious features (Sagawa et al., 2019). For privacy evaluation, we follow the guidelines from Carlini et al. (2022). We train non-overfit models and report the full log-scale ROC curves, the true positive rate (TPR) at a low false positive rate (FPR) region, and also the AUROC curve for completeness. We train 32 shadow models for Waterbirds/CelebA and 16 for FMoW/MultiNLI.
@@
Across all the spurious correlated datasets, the group performance disparity is consistently present, with the spurious groups having the lowest accuracy among all the groups (see Table 5). For example, in Waterbirds, ERM has a test average accuracy of 81.08% while one of the spurious group only 34.41%. Beyond these performance disparity, we now show that spurious correlations also cause privacy issues. Using the state-of-the-art MIA method LiRA (Carlini et al., 2022), we analyze the privacy leakage of each group of the five spurious correlated datasets. For each dataset, we train the shadow models using 50% of the sampled training data as in the LiRA algorithm. We ensure that the sampled subset maintains a similar group proportion as the original dataset by first sampling per group, and then combining all the sampled groups together. Our results in Figure 1 show that across the datasets, there exists a spurious group that exhibits higher privacy leakage than non-spurious groups. The largest disparity is observed at ${\approx}3\%$ FPR area of Waterbirds, where the samples in the most spurious group are ${\approx}10$ times more vulnerable than samples in the non-spurious group. In CelebA, we continue to observe a significant disparity, with the most spurious group being ${\approx}10$ times more vulnerable than the least spurious group. In both the text datasets MultiNLI and CivilComments, the disparity is milder with ${\approx}4$ times difference between the most and least vulnerable groups (see Table 1 for the exact TPR at low FPR). The existence of spurious privacy leakage unfairly exposes some data groups, allowing an attacker to craft better targeted attacks. Our results on spurious correlation extend prior related research on fairness (Zhong et al., 2022; Kulynych et al., 2022; Tian et al., 2024), where they also found privacy disparity between different subpopulations. Surprisingly, we do not observe a spurious privacy leakage in the FMoW dataset, which we investigate in the next section.
@@ 3.2 Connecting privacy leakage with task complexity
Experiment setup. We create two new datasets with 16 (FMoW16) and 4 (FMoW4) classes by sequentially clustering the 62 classes of the original FMoW. We train 16 shadow models for each dataset as in Section 3.1 and use LiRA for privacy analysis. The results are averaged over 5 different target models.
Figure 2(a) shows the decrease of average privacy risk over the total dataset as the task simplifies from 62 to 4 classes (black dot-dashed line). This observation is consistent with prior works on balanced datasets. For example, CIFAR100 has a higher privacy leakage compared to its counterpart CIFAR10 (Shokri et al., 2017; Carlini et al., 2022). The same is found in segmentation tasks where a larger output dimension has higher MIA vulnerability (Shafran et al., 2021). These findings confirm the importance of task complexity on privacy risks. We continue the analysis by investigating at a per-group level on top of the average one. Interestingly, we observe that the privacy disparity between the spurious and non-spurious groups grows as the task complexity simplifies. While the leakage for most of the groups drops, the spurious group 2 remains consistently vulnerable at 6% TPR at 0.1% FPR across the dataset with different classes.
We hypothesize that spurious groups are characterized by similar and fewer discriminative features across tasks of different complexity (e.g. FMoW62 vs FMoW4), contributing to spurious privacy leakage. To quantify the feature similarity, we compute linear centered kernel alignment (CKA; Kornblith et al. (2019)) between feature embeddings of two models. For feature complexity, we apply principal component analysis (PCA) and measure the number of components required to explain a variance threshold $\tau$. Formal definitions are provided in Section B.1. First, we find that CKA scores between spurious group embeddings across FMoW62 and FMoW4 are significantly higher than those for non-spurious groups, indicating more similar feature representations (Figure 2(b); blue bar). Second, PCA analysis show that models trained on the simpler FMoW4 task require far fewer components to reach $\tau=0.95$ compared to FMoW62 (3 vs 25 components), indicating reduced feature complexity (Figure 2(c)). Within each task, spurious groups also require fewer components than non-spurious groups. These results suggest that task simplification leads to increased feature alignment in spurious groups and reduced feature complexity, which exacerbate privacy disparity. In contrast, more complex tasks lead to richer representations but also increase the average privacy vulnerability across groups.

### v2

@@ Abstract
Neural networks trained on real-world data often exhibit biases while simultaneously being vulnerable to privacy attacks aimed at extracting sensitive information. Despite extensive research on each problem individually, their intersection remains poorly understood. In this work, we investigate the privacy impact of spurious correlation bias. We introduce spurious privacy leakage, a phenomenon in which spurious groups are significantly more vulnerable to privacy attacks than non-spurious groups. We observe that privacy disparity between groups increases in tasks with simpler objectives (e.g. fewer classes) due to spurious features. Counterintuitively, we demonstrate that spurious robust methods, designed to reduce spurious bias, fail to mitigate privacy disparity. Our analysis reveals that this occurs because robust methods can reduce reliance on spurious features for prediction, but do not prevent their memorization during training. Finally, we systematically compare the privacy of different model architectures trained with spurious data, demonstrating that, contrary to previous work, architectural choice can affect privacy evaluation.
@@ 1 Introduction
In this work, we examine the privacy consequences of spurious correlation bias, a statistical relationship between two variables that appears to be causal but is either caused by a third confounding variable or randomness. Such correlations are prevalent in real-world datasets, particularly in domains with limited and imbalanced group representation. To address the problem, several mitigation methods have been proposed in the literature with the objective to improve the performance of the worst-group (Sagawa et al., 2019; Nam et al., 2020; Liu et al., 2021a; Izmailov et al., 2022; Yang et al., 2023). However, their privacy implications have not been studied. On the other hand, privacy research typically focuses on simple datasets, such as CIFAR (Krizhevsky, 2009; Hu et al., 2022), overlooking the privacy risks associated with realistic applications that often use biased datasets. These limitations can lead to additional concerns. From an ethical view, an unequal privacy treatment exacerbates existing inequalities, particularly when spurious correlations are sensitive attributes (e.g. race or gender). From a technical perspective, privacy research often assumes that attacks equally affect all samples in the training data. But a spurious correlated dataset can be represented as subgroups with different properties. This raises a natural question: is there a significant privacy risk difference between subgroups? For example, consider a medical dataset where ’young/healthy’ and ’old/sick’ dominate, with a few ’young/sick’ cases. A model can memorize this uncommon (spurious) correlation, at the risk of exposing the ’young/sick’ patients to further privacy vulnerability. This inequality adds on top of the fairness problem: groups already suffering from poor performance can simultaneously face higher privacy risks, violating both performance and privacy equity principles. In such a scenario, how can an auditor detect these fairness and privacy issues related to data biases?
Motivated by these observations, we investigate privacy with spurious correlations to understand how real-world dataset biases can affect privacy vulnerabilities. In practice, we analyze the privacy of deep neural network models trained on real-world spurious correlated datasets using membership inference attacks (MIA), a family of privacy attacks commonly used for their simplicity and versatility (Murakonda & Shokri, 2020; Carlini et al., 2021).
@@
Contributions. We identify a phenomenon we term spurious privacy leakage, where groups with spurious correlation are significantly more vulnerable to MIA than other groups, creating a privacy disparity (Section˜3.1). This phenomenon adds a fairness requirement on top to the existing privacy challenges, modeling a realistic scenario for data-sensitive applications. Prior works demonstrated how privacy auditing may naively conclude that a model satisfies the privacy requirements by evaluating on an aggregated average metric over the entire dataset (Carlini et al., 2022; Ye et al., 2022; Sablayrolles et al., 2019). We demonstrate that this also applies to realistic spurious data, where an average metric evaluation is misleading for evaluating spurious and non-spurious groups. Addressing this privacy disparity is important to bridge and advance the privacy and spurious correlation research, to characterize the risks of the model, and to improve the auditing process.
@@
We continue with further analysis on spurious privacy leakage. Prior works suggested that improving fairness can, in limited scenarios, bound privacy disparity (Kulynych et al., 2022). Given the range of methods proposed to mitigate spurious correlations (equivalent as improving fairness between different data groups) (Sagawa et al., 2019; Kirichenko et al., 2022; Liu et al., 2021a; Zhang et al., 2022), we investigate whether these mitigations can also help to reduce privacy leakage. To explore this, we apply robust training methods to mitigate spurious correlations and re-evaluate the privacy vulnerability. Surprisingly, we find no consistent and significant privacy improvement (Section˜4.1). This observation leads us to introduce a new perspective on group privacy disparity based on memorization (Zhang et al., 2021): spurious robust training improves worst-class performance but does not reduce the memorization level of data compared to standard training (Figure˜8), and thus neither can mitigate the privacy risks. Additional results confirm that robust training essentially learns the same features as standard training. Moreover, we also assess the effectiveness of differential privacy to mitigate the spurious privacy leakage. We observe that given a pretrained model, a stricter privacy guarantee can offer worst group privacy protection at the expense of reduced utility (Section˜4.2).
@@
Lastly, we analyze the influence of model architecture on privacy disparity. While prior works mostly focus on ResNet-like architecture (Carlini et al., 2022; Liu et al., 2022a), we fairly compare the performance and privacy of different model families. Our evaluation include state-of-the-art convolutional and transformer-based architectures, pretrained using both supervised and self-supervised training. Contrary to prior works showing that the best shadow architecture always matches the target one (Carlini et al., 2022; Liu et al., 2022a), our results provide counterexamples when models are trained on spurious correlated data. This highlights a key difference in privacy behavior when models are trained on biased real-world data. We release the code at https://github.com/orientino/spurious-mia.
@@ 2 Background & Related work
where the dataset is divided into $|{\mathcal{G}}|$ groups. The new dataset is $\mathcal{D}=\{(\bm{x}_{i},\bm{y}_{i},\bm{g}_{i})\}$ where $\bm{g}\in{\mathcal{G}}$ is a discrete-valued label (e.g. all the combinations of birds and backgrounds or geographical area information (Koh et al., 2021)). DRO is considered an oracle method due to its explicit use of the group information for the training (Liu et al., 2021a). Additional methods in the literature suppress the spurious features by learning and assigning a different weight per sample (Liu et al., 2021a; Nam et al., 2020), by retraining the classifier head at the end of the training (Kirichenko et al., 2022; Izmailov et al., 2022; Kang et al., 2020), by group sampling (Yang et al., 2024; Idrissi et al., 2022), or using contrastive methods (Zhang et al., 2022).
@@
Privacy and safety. The intersection of privacy and ML safety topics has been extensively studied. Wang et al. (2021) focused on how pruning can mitigate privacy attacks, Shokri et al. (2021) explored the connection between privacy and explainability, Song et al. (2019) found that adversarial training can increase privacy leakage, but however, Li et al. (2024) reported contradictory findings when using a better privacy evaluation (Carlini et al., 2022). In our work, we investigate the privacy risk of real-world spurious correlated datasets, which is related to fairness where the goal is to ensure equal performance across groups. Prior works on privacy-fairness reported that subpopulations can exhibit varying levels of privacy risk (Tian et al., 2024; Zhong et al., 2022; Kulynych et al., 2022). In particular, Tian et al. (2024) showed that fairness methods can mildly mitigate MIA risks using the average metrics. Instead, we use a per-group analysis, revealing the spurious privacy leakage phenomenon with group privacy disparity between spurious and non-spurious data. Kulynych et al. (2022) and Zhong et al. (2022) also investigated privacy disparity on synthetic and tabular datasets. They hypothesize that, under certain conditions, group fairness improvements or differential privacy can be effective as privacy mitigation. In Section˜4.1, we show that neither approaches can effectively address privacy disparities in real-world datasets. Similar to our setting, Yang et al. (2022) studied the privacy risks associated with a toy dataset (MNIST) using a suboptimal evaluation. In contrast, we use state-of-the-art methods to evaluate real-world datasets. Lastly, while out-of-distribution data are known to be more vulnerable to privacy attacks than in-distribution data (Carlini et al., 2022), spurious correlated data differ from out-of-distribution data, as they are by definition in-distribution. Our contributions present a set of results to resolve the conflicts in the literature while also offering new directions for future works.
@@ 3.1 Spurious privacy leakage
Experiment setup. We select the datasets that are used by the spurious correlation community (Yang et al., 2023): Waterbirds (Sagawa et al., 2019), CelebA (Liu et al., 2015), FMoW (Koh et al., 2021), MultiNLI (Williams et al., 2017), and CivilComments (Koh et al., 2021). These datasets contain real-world spurious correlations, diverse modalities, and different target complexity (see Appendix˜A for details). Moreover, to the best of our knowledge, we are the first to study MIA attacks on subgroups of these realistic datasets. We use the pretrained ResNet50 (He et al., 2016) on ImageNet1k and finetune using random crop and horizontal flip. For text datasets, BERT’s bert-base-uncased model (Devlin et al., 2019) is used. We perform hyperparameter optimization for each dataset using a grid search over learning rate (lr), weight decay (wd), and epochs. The grid search and its best hyperparameters are in Appendix˜B. We report the training and test accuracy to evaluate the performance and their difference to quantify the overfitting level. We do the same for the worst-group accuracy (WGA), which is a commonly used proxy metric to measure the mitigation success of spurious features (Sagawa et al., 2019). For privacy evaluation, we follow the guidelines from Carlini et al. (2022). We train non-overfit models and report the full log-scale ROC curves, the true positive rate (TPR) at a low false positive rate (FPR) region, and also the AUROC curve for completeness. We train 32 shadow models for Waterbirds/CelebA and 16 for FMoW/MultiNLI.
@@
Across all the spurious correlated datasets, the group performance disparity is consistently present, with the spurious groups having the lowest accuracy among all the groups (see Table˜5). For example, in Waterbirds, ERM has a test average accuracy of 81.08% while one of the spurious group only 34.41%. Beyond these performance disparity, we now show that spurious correlations also cause privacy issues. Using the state-of-the-art MIA method LiRA (Carlini et al., 2022), we analyze the privacy leakage of each group of the five spurious correlated datasets. For each dataset, we train the shadow models using 50% of the sampled training data as in the LiRA algorithm. We ensure that the sampled subset maintains a similar group proportion as the original dataset by first sampling per group, and then combining all the sampled groups together. Our results in Figure˜1 show that across the datasets, there exists a spurious group that exhibits higher privacy leakage than non-spurious groups. The largest disparity is observed at ${\approx}3\%$ FPR area of Waterbirds, where the samples in the most spurious group are ${\approx}10$ times more vulnerable than samples in the non-spurious group. In CelebA, we continue to observe a significant disparity, with the most spurious group being ${\approx}10$ times more vulnerable than the least spurious group. In both the text datasets MultiNLI and CivilComments, the disparity is milder with ${\approx}4$ times difference between the most and least vulnerable groups (see Table˜1 for the exact TPR at low FPR). The existence of spurious privacy leakage unfairly exposes some data groups, allowing an attacker to craft better targeted attacks. Our results on spurious correlation extend prior related research on fairness (Zhong et al., 2022; Kulynych et al., 2022; Tian et al., 2024), where they also found privacy disparity between different subpopulations. Surprisingly, we do not observe a spurious privacy leakage in the FMoW dataset, which we investigate in the next section.
@@ 3.2 Connecting privacy leakage with task complexity
Experiment setup. We create two new datasets with 16 (FMoW16) and 4 (FMoW4) classes by sequentially clustering the 62 classes of the original FMoW. We train 16 shadow models for each dataset as in Section˜3.1 and use LiRA for privacy analysis. The results are averaged over 5 different target models.
Figure˜2(a) shows the decrease of average privacy risk over the total dataset as the task simplifies from 62 to 4 classes (black dot-dashed line). This observation is consistent with prior works on balanced datasets. For example, CIFAR100 has a higher privacy leakage compared to its counterpart CIFAR10 (Shokri et al., 2017; Carlini et al., 2022). The same is found in segmentation tasks where a larger output dimension has higher MIA vulnerability (Shafran et al., 2021). These findings confirm the importance of task complexity on privacy risks. We continue the analysis by investigating at a per-group level on top of the average one. Interestingly, we observe that the privacy disparity between the spurious and non-spurious groups grows as the task complexity simplifies. While the leakage for most of the groups drops, the spurious group 2 remains consistently vulnerable at 6% TPR at 0.1% FPR across the dataset with different classes.
We hypothesize that spurious groups are characterized by similar and fewer discriminative features across tasks of different complexity (e.g. FMoW62 vs FMoW4), contributing to spurious privacy leakage. To quantify the feature similarity, we compute linear centered kernel alignment (CKA; Kornblith et al. (2019)) between feature embeddings of two models. For feature complexity, we apply principal component analysis (PCA) and measure the number of components required to explain a variance threshold $\tau$. Formal definitions are provided in Section˜B.1. First, we find that CKA scores between spurious group embeddings across FMoW62 and FMoW4 are significantly higher than those for non-spurious groups, indicating more similar feature representations (Figure˜2(b); blue bar). Second, PCA analysis show that models trained on the simpler FMoW4 task require far fewer components to reach $\tau=0.95$ compared to FMoW62 (3 vs 25 components), indicating reduced feature complexity (Figure˜2(c)). Within each task, spurious groups also require fewer components than non-spurious groups. These results suggest that task simplification leads to increased feature alignment in spurious groups and reduced feature complexity, which exacerbate privacy disparity. In contrast, more complex tasks lead to richer representations but also increase the average privacy vulnerability across groups.